Can you help me with this equation problem?
A pasture is twice as long as it is wide. Its area is 115,200 ft^2. How wide is the pasture?

Answer by checkley71(8403)

LENGTH*WIDTH=AREA
2X*X=115,200
2X^2=115,200
X^2=115,200/2
X^2=57,600
X=SQRT57,600
X=240 WHICH IS THE WIDTH
2*240=480 WHICH IS THE LENGTH.
PROOF
240*480=115,200
115,200=115,200